OpenStudy (anonymous):

what are the zeros of f(x)=x^4-7x^3+15x^2-9x?

OpenStudy (anonymous):

To fnd zeros, f(x)=0 x^4-7x^3+15x^2-9x=0

OpenStudy (anonymous):

Now, SOLVE FOR x can u?

OpenStudy (anonymous):

I don't remember can you show me how please

OpenStudy (anonymous):

Try to factor x^4-7x^3+15x^2-9x

OpenStudy (anonymous):

oh okay I remember what to do now!

OpenStudy (anonymous):

Can you show me how to factor that like I remember what to do once you find X but I'm not so great at factoring 4terms and 3terms..

OpenStudy (anonymous):

x^4-7x^3+15x^2-9x =x(x^3 -7x^2+15x-9) =x{ x^3 -x^2 -6x^2 +6x +9x-9} =x{ x^2(x-1) -6x(x-1) +9(x-1)} =x{(x-1)(x^2-6x+9)} =x(x-1){ x^2 -2*x*3+3^2} =x(x-1)(x-3)(x-3)

OpenStudy (anonymous):

Now, can u do the rest?

OpenStudy (anonymous):

Yes! Thank You so much! The X on the outside is equal to? 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yay okay thank you!

OpenStudy (anonymous):

welcome